dayungzy1:

Which of the following statements most accurately describes the difference between 3x^0 and (3x)^0, if a difference exists?

dayungzy1:

don't mind the subject its math lol

umm:

@Ultrilliam - Belongs in Mathematics.

dayungzy1:

duhh I just said don't mind the subject dh

umm:

Lol, I know. He usually put them in the proper subject. Also, welcome to QuestionCove.

jhonyy9:

attention - we know that indifferent every numbers or variables on power of 0 are equal one - so from this result that 3x^0 = 3*1 = 3 end the other term is (3x)^0 = 1 hope is understandably and helped